# Gravity as a deformed topological gauge theory

### v5

@@ 1 Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri gravity and BF theory, see e.g. [1]. They can be mathematically described in terms of Cartan geometries on a manifold $\mathscr{M}$ modeled on a homogeneous space $G/H$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their corresponding Lie algebras. When the geometry is reductive, namely, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$, a Ad$(H)$-invariant decomposition, the associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$, one with values in $\mathfrak{h}$ and the other one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part is arranged to be $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda_{0}}{3\epsilon}}\begin{pmatrix}0&\beta\\ -\epsilon\bar{\beta}&0\end{pmatrix}$ is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$) where $\bar{\beta}=\eta(\beta)$ with $\eta$ the Minkowski metric.
@@
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a $1$-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$ the Lie algebra of $G$ satisfying the conditions [2, p128]:
@@
and $\quad\ \,\text{and }Ad\text{ the adjoint representation of }H\subset G\text{ on the Lie algebra }\mathfrak{g}.$
@@
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$ originating from the $\mathfrak{so}(m)$-valued connection $1$-form $\omega_{\mathfrak{so}}$. This is an $SO(m)$-invariant polynomial.
For a $4$-dimensional manifold one has:
@@
Section 3 will be devoted to the construction of a similar action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group) and $H=CO(3,1)=SO(3,1)\times(\mathbb{R}_{+}\setminus\{0\})$ the Lorentz×$\times$×Weyl group of signature $(3,1)$, so that the infinitesimal Klein geometry $(\mathfrak{g},\mathfrak{h})$ in the sense of [11] turns out to be a symmetric Lie algebra. Then, mutating the $(\mathfrak{g},\mathfrak{h})$ model allows to mimic what we did in the previous Lorentz case. Reducing the quotient $\mathfrak{g}/\mathfrak{h}$ to be of dimension $4$ by constraining the "pair of frames" $(\alpha,\beta)$ one can retrieve an action comprising all terms described in the previous example + a kinetic term for a scalar field (dilations).
@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
with $\bar{\varphi}_{\mathfrak{p}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$ and where $k_{1},k_{2}$ may a priori be considered as 00-form scalar fields. It is worthwhile to notice that this mutation transforms the geometry modeled on $G/H$ given earlier to the Lorentzian Cartan geometry modeled on $G^{\prime}/H$ with elements $g^{\prime}\in G^{\prime}$ obeying the group relation $g^{\prime T}Ng^{\prime}=N$, where $N$ is the following metric:
@@ 2.2 Hodge *-operator
Let $\varepsilon_{a_{1}a_{2}...a_{m}}$ be the Levi-Civita symbol in dimension $m$, such that $\varepsilon_{12...m}=1$. From the metric $g$ and its determinant $|g|$, one can define a Hodge star operator * that acts on $\omega$ as:
@@ 2.3 Associated action
$\hskip-8.53581ptS_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=\int_{\mathscr{M}}\dfrac{36k^{2}\kappa\zeta^{2}}{\ell^{4}}h^{\varepsilon}\bigl{(}i\psi,\eta_{ab}\gamma^{a}\beta^{b}\wedge\tilde{*}D_{\mathfrak{h}}\psi\bigl{)}=\mathrm{Re}S_{D}$ (3)
@@
We can now consider $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{3k}{\Lambda_{0}}$ as fixed. Upon identifying the B-I parameter as $\gamma=\dfrac{e}{r+2y}$, only two degrees of freedom corresponding to the relative value of the parameters $r$ and $y$ and to $|k|$ (the absolute value of $k$) are left.
@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
The geometry modeled on the present $8$-dimensional quotient $G/H$ describes in general an $8$-dimensional manifold $\widetilde{\mathscr{M}}$ invariant under Lorentz transformations (elements in $SO(3,1)$) and dilations (elements of the multiplicative group $\mathbb{R}^{+}$) with soldering form $\varpi_{\mathfrak{m}}$. In the next section, we will see how we can obtain a description of a $4$-dimensional submanifold $\mathscr{M}$ of $\widetilde{\mathscr{M}}$ by imposing a condition on the soldering form (the $\mathfrak{m}$-valued part of the Cartan connection).
@@
If we choose a Cartan geometry such that $\Theta=0$, $f=0$ and Ric$(F)=0$, [11, Chapter 7, Def 2.7] then, as described in [27, p.52], $\bar{\Pi}$ and $\bar{\alpha}$ become, respectively, the Cotton and Schouten tensors, while $F$ would correspond to the Weyl tensor. These tensors constitute the building blocks of conformal gravity.
@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
where $\mathscr{M}_{i}$ is either the $8$-dimensional manifold $\widetilde{\mathscr{M}}$ defined by the Cartan geometry $G/H$ or a $4$-dimensional submanifold $\mathscr{M}\subset\widetilde{\mathscr{M}}$ defined by the identification $\alpha=\dfrac{k}{2}\beta$ [^17] This ought to resemble a gauge fixing which amounts to freezing the degrees of freedom of the $\mathfrak{g}_{1}$-part to be those of the $\mathfrak{g}_{-1}$-sector. It can be linked with [11, Proof of Proposition 3.1, p.285] where the two tetrads are taken to be linearly dependent each other. To some extent, this selects a mutated Weyl geometry within the mutated Möbius one. One thus gets $\dim\mathscr{M}=11-7=4$. One may also consider manifolds of other dimensions by imposing other constraints that reduce the number of degrees of freedom of the soldering form $\varpi_{\mathfrak{m}}$.
@@ 3.2.1 Fermionic matter field action
$\hskip-22.76219ptS_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,\eta_{ab}\dfrac{64k^{2}\kappa^{2}\zeta}{(\ell\ell^{\prime})^{2}}\gamma^{a}\beta^{b}\wedge\tilde{*}D_{\mathfrak{so}}\psi\bigl{)}=\mathrm{Re}S_{D}\,.$ (13)
@@ 3.2.2 The total action
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~{}\mu\nu}\varepsilon^{a~{}\mu\nu}_{~{}b}=0$) as is described in [28, Table 3]. The end result being that the new torsion $T^{a}_{~{}bc}=Z^{a}_{~{}bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of $20$ independent components instead of the usual $24$. Here $Z^{a}_{~{}bc}$ is a $3$-indices tensor and $t_{c}$ is a $4$-dimensional vector with respectively $16$ and $4$ independent components.
@@
$8$-dimensional manifold $\widetilde{\mathscr{M}}$ action
In the case $\alpha\neq\dfrac{k}{2}\beta$ corresponding to the $8$-dimensional manifold $\widetilde{\mathscr{M}}$ ($\alpha$ is related to the special conformal transformations) the action (1) becomes:
@@ 4 Topological total action
This latter relation allows to relax some conditions on torsion of the Cartan connection in order to secure (4) to be 00 ($P(\bar{\Omega}_{\mathfrak{h}}^{n})$ is closed) if one wishes to avoid to consider the $[\mathfrak{m},\mathfrak{m}]$ contributions. Indeed, the Bianchi identity $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ infers that $[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ if the covariant derivative of the torsion $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]=0$. One thus gets
@@
As can be seen in $(a),(b)(c)$ and $(d)$, the vanishing of the $[\mathfrak{m},\mathfrak{m}]$ commutators removes the obstruction to the Chern-Weil theorem in the case of symmetric Lie algebras. Furthermore, as stated in [31, Proposition 4.1] there is an equivalence between a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ with $[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ and an abelian reductive decomposition $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ with $[\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=0$ through a mutation map and where $\mathfrak{m}\simeq\mathfrak{m}^{\prime}$ as $\mathfrak{h}$-modules, (i.e. $[\mathfrak{h},\mathfrak{m}]\subset\mathfrak{m}$ and $[\mathfrak{h},\mathfrak{m}^{\prime}]\subset\mathfrak{m}^{\prime}$ and $\mathfrak{m}\simeq\mathfrak{m}^{\prime}$ as vector spaces, see also [11, Lemma 6.4, p.220]). Exploiting this fact, one has
@@
where we tuned $v$ such that $v,v(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})$ tend towards 00. In this way, we retrieve (if we assume $\mathscr{M}$ has no boundary or when neglecting boundary terms) the equations of motion of section 2 with $\Lambda_{0}\rightarrow 0$. Interestingly, we can make $\Lambda_{0}\rightarrow 0^{\pm}$ such that we end up with a dS or AdS spacetime. To some extent, this limit seems to correspond to the interpretation as a "regulator" of the bare cosmological constant in the case of AdS spacetime [20].
@@ 5 Conclusion
The second example deals with the Möbius geometry. In this case restricting the geometry to a $4$-dimensional submanifold $\mathscr{M}$ yields the same results as in the previous class of examples if one drops out dilations. Adding to the original action (3.2) an interaction $T^{a}\wedge\lambda\wedge\beta_{a}$ between dilations and torsion instead leads to Einstein’s equations (modified by the Holst term) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $-\dfrac{2\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.
@@
Acknowledgments
We thank our colleagues Thomas Schücker, Pietro Dona and Federico Piazza for helpful discussions. One of us (J.T) would also like to thank Thomas Ströbl for insightful comments.

### v6

@@ 1 Introduction
[…]
@@
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a 111-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$ the Lie algebra of $G$ satisfying the conditions [2, p128]:
@@
$\quad\ \,\text{and }Ad\text{ the adjoint representation of }H\subset G\text{ on the Lie algebra }\mathfrak{g}.$
@@
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$ originating from the $\mathfrak{so}(m)$-valued connection 111-form $\omega_{\mathfrak{so}}$. This is an $SO(m)$-invariant polynomial.
For a 444-dimensional manifold one has:
@@
Section 3 will be devoted to the construction of a similar action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group) and $H=CO(3,1)=SO(3,1)\times(\mathbb{R}_{+}\setminus\{0\})$ the Lorentz×$\times$×Weyl group of signature (3,1)(3,1)( 3 , 1 ), so that the infinitesimal Klein geometry $(\mathfrak{g},\mathfrak{h})$ in the sense of [11] turns out to be a symmetric Lie algebra. Then, mutating the $(\mathfrak{g},\mathfrak{h})$ model allows to mimic what we did in the previous Lorentz case. Reducing the quotient $\mathfrak{g}/\mathfrak{h}$ to be of dimension 444 by constraining the "pair of frames" $(\alpha,\beta)$ one can retrieve an action comprising all terms described in the previous example + a kinetic term for a scalar field (dilations).
@@ 2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
with $\bar{\varphi}_{\mathfrak{p}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$ and where $k_{1},k_{2}$ may a priori be considered as 0-form scalar fields. It is worthwhile to notice that this mutation transforms the geometry modeled on $G/H$ given earlier to the Lorentzian Cartan geometry modeled on $G^{\prime}/H$ with elements $g^{\prime}\in G^{\prime}$ obeying the group relation $g^{\prime T}Ng^{\prime}=N$, where $N$ is the following metric:
@@ 2.2 Hodge *-operator
[…]
@@ 2.3 Associated action
$\thinspace S_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=\int_{\mathscr{M}}\dfrac{36k^{2}\kappa\zeta^{2}}{\ell^{4}}h^{\varepsilon}\bigl{(}i\psi,\eta_{ab}\gamma^{a}\beta^{b}\wedge\tilde{*}D_{\mathfrak{h}}\psi\bigl{)}=\mathrm{Re}S_{D}$ (3)
@@
[…]
@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
The geometry modeled on the present 888-dimensional quotient $G/H$ describes in general an 888-dimensional manifold $\widetilde{\mathscr{M}}$ invariant under Lorentz transformations (elements in $SO(3,1)$) and dilations (elements of the multiplicative group $\mathbb{R}^{+}$) with soldering form $\varpi_{\mathfrak{m}}$. In the next section, we will see how we can obtain a description of a 444-dimensional submanifold $\mathscr{M}$ of $\widetilde{\mathscr{M}}$ by imposing a condition on the soldering form (the $\mathfrak{m}$-valued part of the Cartan connection).
@@
[…]
@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
where $\mathscr{M}_{i}$ is either the 888-dimensional manifold $\widetilde{\mathscr{M}}$ defined by the Cartan geometry $G/H$ or a 444-dimensional submanifold $\mathscr{M}\subset\widetilde{\mathscr{M}}$ defined by the distribution:
Proposition 3.1.
The subset of vector fields $\mathscr{D}_{k}=\{X\in\Gamma(T\widetilde{\mathscr{M}})|\alpha(X)=\dfrac{k}{2}\beta(X)\}$ defines an integrable submanifold of dimension four $\mathscr{M}\subset\widetilde{\mathscr{M}}$.
Proof.
Indeed, $\mathscr{D}_{k}=\{X\in\Gamma(T\widetilde{\mathscr{M}})|\alpha(X)=\dfrac{k}{2}\beta(X)\}$ provides us with the vielbein $\alpha-\dfrac{k}{2}\beta$ which corresponds to the four Pfaff forms defining the submanifold $\mathscr{M}$ such that $\alpha-\dfrac{k}{2}\beta=0$ on $\mathscr{M}$. For all $X,Y\in\mathscr{D}_{k}$ one has $(\Pi-\lambda\wedge\alpha)(X,Y)=\dfrac{k}{2}(\Theta+\lambda\wedge\beta)(X,Y)=\dfrac{k}{2}T(X,Y)$ which yields
$d(\alpha-\dfrac{k}{2}\beta)=-A\wedge(\alpha-\dfrac{k}{2}\beta).$
Hence, by the Frobenius theorem [11, Prop.5.3, p.81] the corresponding distribution defines $\mathscr{M}$ as an integrable submanifold of $\widetilde{\mathscr{M}}$. Therefore, there is a subclass of conformal Cartan connections characterizing $\mathscr{M}$; this subclass can be reached through this kind of "gauge fixing".171717 We are very indebted to Thierry Masson for his key input in the construction of this integrable distribution defined by this specific Cartan connection. This ought to resemble a gauge fixing which amounts to freezing the degrees of freedom of the $\mathfrak{g}_{1}$-part to be those of the $\mathfrak{g}_{-1}$-sector. It can be linked with [11, Proof of Proposition 3.1, p.285] where the two tetrads are taken to be linearly dependent each other. To some extent, this selects a mutated Weyl geometry within the mutated Möbius one. One thus gets $\dim\mathscr{M}=11-7=4$. One may also consider manifolds of other dimensions by imposing other constraints that reduce the number of degrees of freedom of the soldering form $\varpi_{\mathfrak{m}}$.
@@ 3.2.1 Fermionic matter field action
$\thinspace S_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,\eta_{ab}\dfrac{64k^{2}\kappa^{2}\zeta}{(\ell\ell^{\prime})^{2}}\gamma^{a}\beta^{b}\wedge\tilde{*}D_{\mathfrak{so}}\psi\bigl{)}=\mathrm{Re}S_{D}\,.$ (13)
@@ 3.2.2 The total action
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~\mu\nu}\varepsilon^{a~\mu\nu}_{~b}=0$) as is described in [28, Table 3]. The end result being that the new torsion $T^{a}_{~bc}=Z^{a}_{~bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of 202020 independent components instead of the usual 242424. Here $Z^{a}_{~bc}$ is a 333-indices tensor and $t_{c}$ is a 444-dimensional vector with respectively 161616 and 444 independent components.
@@
888-dimensional manifold $\widetilde{\mathscr{M}}$ action
In the case $\alpha\neq\dfrac{k}{2}\beta$ corresponding to the 888-dimensional manifold $\widetilde{\mathscr{M}}$ ($\alpha$ is related to the special conformal transformations) the action (1) becomes:
@@ 4 Topological total action
This latter relation allows to relax some conditions on torsion of the Cartan connection in order to secure (4) to be 0 ($P(\bar{\Omega}_{\mathfrak{h}}^{n})$ is closed) if one wishes to avoid to consider the $[\mathfrak{m},\mathfrak{m}]$ contributions. Indeed, the Bianchi identity $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ infers that $[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ if the covariant derivative of the torsion $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]=0$. One thus gets
@@
[…]
@@
where we tuned $v$ such that $v,v(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})$ tend towards 0. In this way, we retrieve (if we assume $\mathscr{M}$ has no boundary or when neglecting boundary terms) the equations of motion of section 2 with $\Lambda_{0}\rightarrow 0$. Interestingly, we can make $\Lambda_{0}\rightarrow 0^{\pm}$ such that we end up with a dS or AdS spacetime. To some extent, this limit seems to correspond to the interpretation as a "regulator" of the bare cosmological constant in the case of AdS spacetime [20].
@@ 5 Conclusion
The second example deals with the Möbius geometry. In this case restricting the geometry to a 444-dimensional submanifold $\mathscr{M}$ yields the same results as in the previous class of examples if one drops out dilations. Adding to the original action (3.2) an interaction $T^{a}\wedge\lambda\wedge\beta_{a}$ between dilations and torsion instead leads to Einstein’s equations (modified by the Holst term) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $-\dfrac{2\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.
@@
Acknowledgements
We thank our colleagues Thierry Masson, Thomas Schücker, Pietro Dona and Federico Piazza for helpful discussions. One of us (J.T) would also like to thank Thomas Ströbl for insightful comments.